A newly synthesized drug (molecular mass = 322 g/mol) was tested for its optical activity. To do this, 15.0 grams of the drug was dissolved in water and brought to a volume of 12 mL. This solution was placed in a 7.50 cm cell and the optical rotation was found to be -31.0°. What is the specific rotation for this new drug?
Answer #1

and V= 12 -12cm length cl tube-75-cm Mass ot dowg-15.0 31 2 -3.306. cm- ㄒㄧ 123 .5 cm) 3.304c Because Nagtiue and posi tiv sign ndiate clockuoise 3tespe clivelyor

Number of moles of drug = (15g)/(322g/mol) = 0.04658mol

Concentration of drug = (number of moles of drug)1000÷(volume of solution)

Concentration of drug = (0.04658mol)(1000/12L) = 3.881mol/L = 3.881M